# xCOMET-lite: Bridging the Gap Between Efficiency and Quality in Learned MT Evaluation Metrics

### v1

@@ Abstract
State-of-the-art trainable machine translation evaluation metrics like xCOMET achieve high correlation with human judgment but rely on large encoders (up to 10.7B parameters), making them computationally expensive and inaccessible to researchers with limited resources. To address this issue, we investigate whether the knowledge stored in these large encoders can be compressed while maintaining quality. We employ distillation, quantization, and pruning techniques to create efficient xCOMET alternatives and introduce a novel data collection pipeline for efficient black-box distillation. Our experiments show that, using quantization, xCOMET can be compressed up to three times with no quality degradation. Additionally, through distillation, we create an xCOMET-lite metric, which has only 2.6% of xCOMET-XXL parameters, but retains 92.1% of its quality. Besides, it surpasses strong small-scale metrics like COMET-22 and BLEURT-20 on the WMT22 metrics challenge dataset by 6.4%, despite using 50% fewer parameters. All code, dataset, and models are available online.
@@
Daniil Larionov1footnotemark: , Mikhail Seleznyov2, Vasiliy Viskov2, Alexander Panchenko2, Steffen Eger1 1 NLLG, University of Mannheim 2 Skoltech
disable]SE: we can fix and rewrite the abstract in the end. Also promise code+data
disable]SE: which model is xComet-Lite? The distilled? All of them? Can we prune and quantize on top of the distilled? (or vice versa?); DL: the idea is that xCOMET-lite is our distilled model, i’m working now to do some interaction analysis
disable]SE: instead of xComet-Lite, we could use a diminutive from another language, similar to inho in COMETinho. https://en.wikipedia.org/wiki/List_of_diminutives_by_language. German would be "lein". What would Russian be?; DL: i guess something like xCOMETik, but sounds a bit cringe, which could harm disseminations
disable]SE: I am sure the reviewers will complain that there are only 3 language pairs involved. not very much. What about those from Eval4NLP?; DL: eval4nlp covers en-de, en-es and en-zh, adding spanish into the total list of covered languages
00footnotetext: Corresponding Author (daniil.larionov@uni-mannheim.de)
@@ 1 Introduction
disable]SE: in general or in our experiments? I could imagine others have explored this in their tasks earlier; DL: reformulated
disable]SE: fig. 1 is too small. Also, you could cutoff the y-axis below 0.2 so that differences become more clearly visible. Also, you should give the model name "xComet-lite" in Fig. 1. Further, why do you have only a name for the distilled version but not any other version (quantized, pruning)?; DL: re further. that is because distillation has taken most of the efforts (choosing encoder, dataset curation, training).
@@ 2 Related Work
Recent work has explored improving the transparency and capabilities of MT evaluation metrics. Juraska et al. (2023) introduced MetricX. This learned regression-based metric achieves state-of-the-art correlations with human judgments through multi-stage fine-tuning on direct assessment data, consolidated MQM scores, and small-scale synthetic corpora, which is used to boost robustness. It is based on the mT5-XXL encoder-decoder model with 11B parameters. Kocmi and Federmann (2023) proposed GEMBA-MQM, which leverages the GPT-4 language model with a few-shot prompting approach to identify translation error spans and categories. This enables detailed error analysis, though reliance on the computationally expensive proprietary GPT-4 LLM poses challenges for academic research. Guerreiro et al. (2023) developed xCOMET, a learned metric based on the XLM-RoBERTa-XL/XXL encoder that bridges sentence-level quality prediction with fine-grained error span detection. By training on direct assessment and MQM data, xCOMET achieves top quality on sentence-level, system-level, and error span prediction tasks while providing interpretability through its predicted error spans.
@@
Previously, researchers have also explored techniques for creating more efficient MT evaluation metrics while preserving their correlation with human judgments. Kamal Eddine et al. (2022) proposed FrugalScore, which learns lightweight versions of metrics like BERTScore and MoverScore using knowledge distillation. Their distilled metrics perform similarly to the originals while being much faster and having orders of magnitude fewer parameters. Rei et al. (2022b) introduced COMETINHO, a more compact and faster version of the COMET metric. They optimize the COMET code using caching and length batching and further compress the model using pruning and knowledge distillation on synthetic data. The resulting model is 80% smaller and over 2 times faster than the original while maintaining competitive quality. These works studied knowledge distillation and pruning in application to rather simple and small-scale regression metrics like BERTScore, MoverScore, and COMET. In contrast, we scale and adapt these methods for the much larger state-of-the-art xCOMET metric, which generates error annotations in addition to the quality scores. Further, we apply a more comprehensive pruning approach than the one applied in Rei et al. (2022b) as, after pruning layers, we do a fine-tuning of a small selection of model parameters, which substantially reduces performance drop. Additionally, we are the first to study quantization for MT evaluation metrics, as well as do an interaction analysis between compression methods. disable]SE: lines that end in one word can be eliminated by paraphrasing or elimination. That saves space for valuable content
@@
Larionov et al. (2023) provide a comprehensive evaluation of techniques for improving the efficiency of MT evaluation metrics. They explore replacing computation-heavy transformers with more lightweight variants, employing linear and quadratic approximations for alignment algorithms, and using adapters for parameter-efficient training of the COMET metric. Their experiments across multiple datasets and metrics demonstrate that distilled models like TinyBERT can provide an optimal balance of quality and efficiency, with the parallel adapter configuration yielding the best results for the COMET metric training. Unlike this paper, we study compression methods that help boost inference-time efficiency for learned metrics, which helps reduce compute costs and environmental footprint on a much broader scale.disable]SE: and how do we differ?
@@ 3 Methods
disable]SE: by using paragraph instead of subsection, you can save valuable space for real content; DL: changed to paragraphsQuantization is a highly effective compression method with two main approaches: quantization-aware training (QAT) and post-training quantization (PTQ) (Nagel et al., 2021). QAT offers better disable]SE: better? prediction quality but requires costly training, making PTQ more popular. PTQ is further divided into data-free and data-aware methods, where the latter relies on calibration to estimate thedisable]SE:? data distribution parameters for higher prediction quality. Another distinction is weight-only quantization and weight & activation quantization, with the second approach having slightly lower prediction quality but potential for faster computations using efficient 8- or 4-bit CUDA kernels.
@@
Inspired by recent works on layer pruning in LLMs (Gromov et al., 2024; Men et al., 2024) which remove 25-50% of layers with moderate quality drop, we test its applicability for inducing efficient metrics. Specifically, we adopt a simple pruning technique, described in Sec. 4.4 disable]SE: fix hard-coded reference; DL: how can we reference a section in another paper withourd hardcoding it? of Gromov et al. (2024): in an $L$-layer model, we drop layers $(L-n)$ to $(L-1)$. This heuristic is based on the observations that pruning deeper layers should affect the model less, as fewer layers rely on changes made by this layer, but also that the ultimate layer is especially important as it “decodes” the hidden states for the last fragment of the network, and cannot be removed. To mitigate the quality drop incurred by layer removal, we apply parameter-efficient fine-tuning. Concretely, we fine-tune all biases in linear layers, LayerNorm affine parameters, layerwise attention weights, and the regression and tagging heads of xCOMET. This is akin to the BitFit (Zaken et al., 2022) sparse-fine-tuning approach, and has the benefit of adding no parameters and being extremely simple to implement.
@@ 4.1 Evaluation
To evaluate the computational efficiency of compressed models, we measure inference speed in samples per second (samples/s). disable]SE: what about other measures, e.g., memory usage?; DL: added below For a given language pair, we divide the amount of examples by the total time needed to inference the model on the set. Due to the GPU execution and memory models, some operations, such as matrix multiplication, take the same time to execute regardless of the amount of data supplied. Thus, using the largest possible batch size that fits into the accelerator memory is most efficient. To select the optimal batch size, we start with batch size $1$ and increase it by a factor of $2$ until we reach the memory limit on the given GPU. We test model throughput on RTX 3090 and A100 to explore performance on consumer- and production-level GPUs. Additionally, we provide peak vRAM usage for each model on a fixed batch size of 8.
@@ 4.2 Setup
We use the GPTQ (Frantar et al., 2023) quantization algorithm and quantize xCOMET to 8, 4, 3, and 2 bits per parameter. We keep default hyperparameters, including using a small subsample of the C4 (Chung et al., 2022) dataset for calibration. In addition to that, we experiment with data-free quantization methods: LLM.int8() – 8 bit and QLoRA – 4 bit. We use the implementation from the bitsandbytes python library. Initial experiments indicated that models worked faster with their 4-bit quantization implementation if weights were converted to mixed precision beforehand.This observation was also true for 8-bit quantization, but in this case the quality drop became substantial. Thus, we report LLM.int8() without any uncompressed model transformations, and QLoRA with half-precision model weight conversion.
@@
We use mDeBERTa v3 (He et al., 2023) as a student. It has 278 M disable]SE: here and throughout: 278m?; DL: why small m? in literature i always see large M for millions of parameters parameters — 13 times fewer than xCOMET-XL, 39 times fewer than xCOMET-XXL, and 2 times fewer than COMET-22 — one of the top performers in WMT22 Metrics Shared Task. This model was chosen as it shows superior quality on multilingual language understanding tasks such as XNLI (Conneau et al., 2018), compared to alternatives of similar size: InfoXLM (Chi et al., 2021) and XLM-RoBERTa (Conneau et al., 2020). We trained for 1 epoch, with learning rate of $52e-$ 5 for scoring head and $51e-$ 5 for encoder. We set the batch size to 64. Scoring head was configured with two hidden fully connected layers with sizes 3072 and 1024. We compare the prediction quality of the distilled model with original models xCOMET-XL/XXL, as well as with best-performing models of similar size: BLEURT-20 (Sellam et al., 2020) with 579 M parameters and COMET-22 (Rei et al., 2022a) with 581 M parameters.
@@ 4.3 Results
Table 1: Summary table for different compression methods in setting with reference translations. Average Kendall correlation is computed over three language pairs. Underlined values indicate compression methods with best prediction quality. For peak memory the mean and maximum values are computed, and the maximum is reported in parentheses. XL stands for xCOMET-XL, XXL stands for xCOMET-XXL.
@@
Table 3: Speed results for different methods in settings with reference. Importantly, here the memory consumption is higher than in Table 1, as we aim for maximal throughput on a given GPU. Average Kendall correlation is computed over three language pairs. Samples per second are reported for both 3090 and A100 GPUs. XL stands for xCOMET-XL, XXL stands for xCOMET-XXL.
@@
disable]SE: Table 1 needs to say: pruning & quantization, memory vs. quality. Table 3 needs to say in the caption: pruning & quantization throughput vs. quality; DL: But there is also distillation, it is even separated from the rest. SE: ok
@@
Quantization proves highly effective in reducing memory consumption while maintaining quality. For xCOMET-XL, GPTQ 8-bit achieves nearly identical quality to the baseline, with an average Kendall correlation of 0.420, while reducing peak memory usage by 33% (from 7.76 GB to 5.20 GB). GPTQ 3-bit provides the largest memory reduction of 54% (3.54 GB) at the cost of a 0.013 decrease in correlation. Notably, xCOMET-XXL sees no quality degradation with GPTQ 8-bit and 3-bit, despite memory reductions of 38% (from 22.27 GB to 13.81 GB) and 64% (to 7.99 GB), resp. Relative to GPTQ 8- and 3-bit precision, LLM.int8() and QLoRA 4-bit offer memory savings of 7% (7.24 GB) and 27% (5.68 GB), with a slight decrease in correlation (0.020 and 0.012 lower).
In throughput, quantization mostly results in a trade-off between speed and memory consumption. As we see in Table 3, GPTQ 8-bit processes about 53-55% fewer samples/s (10.8-13.9 vs. 23.1-30.9) than the uncompressed xCOMET-XL model on an RTX 3090 GPU. However, LLM.int8() and QLoRA demonstrate increase of throughput on an A100 GPU. The latter method offers 6% acceleration for xCOMET-XL, processing 62.9 samples/s in median terms, while uncompressed model allows to handle 59.5 samples/s. For RTX 3090, both methods demonstrate comparable inference speed.
@@
Pruning layers offers a balance between memory reduction and prediction quality. As shown in Table 1, for xCOMET-XL, pruning 8 and 16 layers reduces memory consumption by 18% (from 7.76 GB to 6.34 GB) and 37% (to 4.90 GB), resp., while incurring a manageable decrease in correlation (0.032 and 0.056 lower). xCOMET-XXL exhibits a similar trend, with pruning 8 and 16 layers reducing memory by 13% (from 22.27 GB to 19.39 GB) and 29% (to 15.91 GB), and lowering correlation by 0.016 and 0.035, resp.
Pruning substantially improves throughput, particularly for xCOMET-XL. As we can see in Table 3, pruning 16 layers nearly doubles the samples processed per second (38.6-51.6 vs. 23.1-30.9, a 67-94% increase) compared to the uncompressed model on an RTX 3090.
@@ 4.4 Interaction Analysis
We can also apply quantization to the distilled model. Unfortunately, due to architectural details, GPTQ quantization is incompatible with the mDeBERTa architecture. Instead, we apply LLM.int8() and QLoRA quantization (8-bit and 4-bit, respectively). When comparing the 8-bit quantized xCOMET-lite model to the non-quantized one, we observe only a marginal drop in correlation with human judgment. The 8-bit model achieves an average score of 0.369 across language pairs with references, compared to original xCOMET-lite 0.388. For pairs without references, the 8-bit model scores 0.354, while xCOMET-lite achieves 0.363. Notably, the model quantized into 4-bit mode yields a slightly higher correlation for pairs with references, namely 0.379. Furthermore, quantization substantially reduces memory usage. The 8-bit quantization decreases the peak memory consumption of the distilled model by 17% from 1.8 GB to 1.5 GB, while the 4-bit quantization further reduces it to 1.4 GB. These results demonstrate that quantization is a viable option for further compressing the distilled model without substantial quality degradation. See Table 5 in Appendix D for full results. disable]SE: So what do these results mean? Are they good / bad? What should I do with numbers out of context?

### v2

@@ Abstract
State-of-the-art trainable machine translation evaluation metrics like xCOMET achieve high correlation with human judgment but rely on large encoders (up to 10.7B parameters), making them computationally expensive and inaccessible to researchers with limited resources. To address this issue, we investigate whether the knowledge stored in these large encoders can be compressed while maintaining quality. We employ distillation, quantization, and pruning techniques to create efficient xCOMET alternatives and introduce a novel data collection pipeline for efficient black-box distillation. Our experiments show that, using quantization, xCOMET can be compressed up to three times with no quality degradation. Additionally, through distillation, we create an 278M-sized xCOMET-lite metric, which has only 2.6% of xCOMET-XXL parameters, but retains 92.1% of its quality. Besides, it surpasses strong small-scale metrics like COMET-22 and BLEURT-20 on the WMT22 metrics challenge dataset by 6.4%, despite using 50% fewer parameters. All code, dataset, and models are available online.
@@
Daniil Larionov1  Mikhail Seleznyov4,3  Vasiliy Viskov3 Alexander Panchenko3,4Steffen Eger1,2 1 NLLG, University of Mannheim, 2 University of Technology Nuremberg, 3 Skoltech, 4 AIRI daniil.larionov@uni-mannheim.de
@@ 2 Related Work
Recent work has explored improving the transparency and capabilities of MT evaluation metrics. Juraska et al. (2023) introduced MetricX. This learned regression-based metric achieves state-of-the-art correlations with human judgments through multi-stage fine-tuning on direct assessment data, consolidated MQM scores, and small-scale synthetic corpora, which is used to boost robustness. It is based on the mT5-XXL encoder-decoder model with 11B parameters. Kocmi and Federmann (2023) proposed GEMBA-MQM, which leverages the GPT-4 language model with a few-shot prompting approach to identify translation error spans and categories.
This enables detailed error analysis, though reliance on the computationally expensive proprietary GPT-4 LLM poses challenges for academic research. Guerreiro et al. (2023) developed xCOMET, a learned metric based on the XLM-RoBERTa-XL/XXL encoder that bridges sentence-level quality prediction with fine-grained error span detection. By training on direct assessment and MQM data, xCOMET achieves top quality on sentence-level, system-level, and error span prediction tasks while providing interpretability through its predicted error spans.
@@
Previously, researchers have also explored techniques for creating more efficient MT evaluation metrics while preserving their correlation with human judgments. Kamal Eddine et al. (2022) proposed FrugalScore, which learns lightweight versions of metrics like BERTScore and MoverScore using knowledge distillation. Their distilled metrics perform similarly to the originals while being much faster and having orders of magnitude fewer parameters. Rei et al. (2022b) introduced COMETINHO, a more compact and faster version of the COMET metric. They optimize the COMET code using caching and length batching and further compress the model using pruning and knowledge distillation on synthetic data. The resulting model is 80% smaller and over 2 times faster than the original while maintaining competitive quality. These works studied knowledge distillation and pruning in application to rather simple and small-scale regression metrics like BERTScore, MoverScore, and COMET.
In contrast, we scale and adapt these methods for the much larger state-of-the-art xCOMET metric, which generates error annotations in addition to the quality scores. Further, we apply a more comprehensive pruning approach than the one applied in Rei et al. (2022b) as, after pruning layers, we do a fine-tuning of a small selection of model parameters, which substantially reduces performance drop. Additionally, we are the first to study quantization for MT evaluation metrics, as well as do an interaction analysis between compression methods.
@@
Larionov et al. (2023) provide a comprehensive evaluation of techniques for improving the efficiency of MT evaluation metrics. They explore replacing computation-heavy transformers with more lightweight variants, employing linear and quadratic approximations for alignment algorithms, and using adapters for parameter-efficient training of the COMET metric. Their experiments across multiple datasets and metrics demonstrate that distilled models like TinyBERT can provide an optimal balance of quality and efficiency, with the parallel adapter configuration yielding the best results for the COMET metric training. Unlike this paper, we study compression methods that help boost inference-time efficiency for learned metrics, which helps reduce compute costs and environmental footprint on a much broader scale.
@@ 3 Methods
Quantization is a highly effective compression method with two main approaches: quantization-aware training (QAT) and post-training quantization (PTQ) (Nagel et al., 2021). QAT offers better prediction quality but requires costly training, making PTQ more popular. PTQ is further divided into data-free and data-aware methods, where the latter relies on calibration to estimate the data distribution parameters for higher prediction quality. Another distinction is weight-only quantization and weight & activation quantization, with the second approach having slightly lower prediction quality but potential for faster computations using efficient 8- or 4-bit CUDA kernels.
@@
Inspired by recent works on layer pruning in LLMs (Gromov et al., 2024; Men et al., 2024) which remove 25-50% of layers with moderate quality drop, we test its applicability for inducing efficient metrics. Specifically, we adopt a simple pruning technique, described in Sec. 4.4 of Gromov et al. (2024): in an $L$-layer model, we drop layers $(L-n)$ to $(L-1)$. This heuristic is based on the observations that pruning deeper layers should affect the model less, as fewer layers rely on changes made by this layer, but also that the ultimate layer is especially important as it “decodes” the hidden states for the last fragment of the network, and cannot be removed. To mitigate the quality drop incurred by layer removal, we apply parameter-efficient fine-tuning. Concretely, we fine-tune all biases in linear layers, LayerNorm affine parameters, layerwise attention weights, and the regression and tagging heads of xCOMET. This is akin to the BitFit (Zaken et al., 2022) sparse-fine-tuning approach, and has the benefit of adding no parameters and being extremely simple to implement.
We also evaluate magnitude pruning and Wanda pruning (Sun et al., 2024). In magnitude pruning, the importance of each weight $S_{ij}$ is directly estimated by its magnitude $|W_{ij}|$. Wanda pruning refines this approach by weighting each $|W_{ij}|$ by the average L2 norm of its corresponding input features, $\frac{1}{N}\sum_{j=1}^{N}\|x_{j}\|_{2}$, aiming to provide a more informed measure of importance. In both methods, the weights with the lowest importance scores are pruned according to the specified sparsity pattern (unstructured, 2:4 or 4:8).
@@ 4.1 Evaluation
To evaluate the computational efficiency of compressed models, we measure inference speed in samples per second (samples/s). For a given language pair, we divide the amount of examples by the total time needed to inference the model on the set. Due to the GPU execution and memory models, some operations, such as matrix multiplication, take the same time to execute regardless of the amount of data supplied. Thus, using the largest possible batch size that fits into the accelerator memory is most efficient. To select the optimal batch size, we start with batch size $1$ and increase it by a factor of $2$ until we reach the memory limit on the given GPU. We test model throughput on RTX 3090 and A100 to explore performance on consumer- and production-level GPUs. Additionally, we provide peak vRAM usage for each model on a fixed batch size of 8.
@@ 4.2 Setup
We use the GPTQ (Frantar et al., 2023) quantization algorithm and quantize xCOMET to 8, 4, 3, and 2 bits per parameter. We keep default hyperparameters, except using a small subsample of the WikiText2 (Merity et al., 2017) dataset for calibration. In addition to that, we experiment with data-free quantization methods: LLM.int8() – 8 bit and QLoRA – 4 bit. We use the implementation from the bitsandbytes python library.Initial experiments indicated that models worked faster with their 4-bit quantization implementation if weights were converted to mixed precision beforehand. This observation was also true for 8-bit quantization, but in this case the quality drop became substantial. Thus, we report LLM.int8() without any uncompressed model transformations, and QLoRA with half-precision model weight conversion.
@@
With Wanda pruning we try 2:4 and 4:8 patterns, to explore setups which can realistically provide speedups on GPU. We use 256 calibration samples from WikiText2333In (Sun et al., 2024) authors use 128 calibration samples from C4, but we couldn’t reproduce the code related to sampling examples from C4., and do not finetune pruned model, as the original method does not require it. We also run simple magnitude pruning with 2:4 and 4:8 sparsity patterns.
@@
We use mDeBERTa v3 (He et al., 2023) as a student. It has 278 M parameters — 13 times fewer than xCOMET-XL, 39 times fewer than xCOMET-XXL, and 2 times fewer than COMET-22 — one of the top performers in WMT22 Metrics Shared Task. This model was chosen as it shows superior quality on multilingual language understanding tasks such as XNLI (Conneau et al., 2018), compared to alternatives of similar size: InfoXLM (Chi et al., 2021) and XLM-RoBERTa (Conneau et al., 2020). We trained for 1 epoch, with learning rate of $52e-$ 5 for scoring head and $51e-$ 5 for encoder. We set the batch size to 64. Scoring head was configured with two hidden fully connected layers with sizes 3072 and 1024. We compare the prediction quality of the distilled model with original models xCOMET-XL/XXL, as well as with best-performing models of similar size: BLEURT-20 (Sellam et al., 2020) with 579 M parameters and COMET-22 (Rei et al., 2022a) with 581 M parameters.
@@ 4.3 Results
Model Compression method Average Kendall correlation Peak memory consumption (GB) mean (max)
@@
Table 1: An overview table with quality / peak memory consumption tradeoff for various representative compression methods in setting with reference translations. Average Kendall correlation and mean/max memory consumption is computed over three language pairs. Underlined values indicate compression methods with best prediction quality. XL stands for xCOMET-XL, XXL stands for xCOMET-XXL. For Wanda pruning, VRAM consumption is reported using the official method implementation, which stores pruned weights as zeros in original precision. However, potentially 4:8 pruning could deliver almost x2 memory usage reduction.
@@
Table 3: Speed results for various methods in settings with reference. Importantly, here the memory consumption is higher than in Table 1, as we aim for maximal throughput on a given GPU. Average Kendall correlation is computed over three language pairs. Samples per second are reported for both 3090 and A100 GPUs. XL stands for xCOMET-XL, XXL stands for xCOMET-XXL. OOM means Out Of Memory error.
@@
Quantization proves highly effective in reducing memory consumption while maintaining quality. For xCOMET-XL, GPTQ 8-bit achieves nearly identical quality to the baseline, with an average Kendall correlation of 0.420, while reducing peak memory usage by 33%. GPTQ 3-bit provides the largest memory reduction of 54% at the cost of a 0.013 decrease in correlation. Notably, xCOMET-XXL sees no quality degradation with GPTQ 8-bit and 3-bit, despite memory reductions of 38% and 64%, respectively. LLM.int8() and QLoRa are suboptimal in terms of quality / peak memory consumption tradeoff, dominated by GPTQ 8-bit and GPTQ 3-bit respectvely.
However, as we see in Table 3, GPTQ slows models down, most likely due to usage of non-optimized CUDA kernels, while QLoRa maintains the thoughput on par with non-compressed model.
@@
Layer pruning substantially improves throughput, particularly for xCOMET-XL. As we can see in Table 3, pruning 16 layers provides  67% speedup compared to the uncompressed model on an RTX 3090. However, the quality drop is larger compared to quantization methods.
Interestingly, magnitude pruning slightly outperforms Wanda pruning, though the latter uses more involved weight importance estimation. Moreover, magnitude pruning performs on par with removing 8 layers, despite keeping only 50% of non-zero weights. Due to some inefficiencies in official implementation, Wanda pruning and magnitude pruning get OOM error on RTX 3090 on some of the datasets; however, we expect they would show speedups similar to ones on A100.
@@ 4.4 Interaction Analysis
We can also apply quantization to the distilled model. Unfortunately, due to architectural details, GPTQ quantization is incompatible with the mDeBERTa architecture. Instead, we apply LLM.int8() and QLoRA quantization (8-bit and 4-bit, respectively). When comparing the 8-bit quantized xCOMET-lite model to the non-quantized one, we observe only a marginal drop in correlation with human judgment. The 8-bit model achieves an average score of 0.369 across language pairs with references, compared to original xCOMET-lite 0.388. For pairs without references, the 8-bit model scores 0.354, while xCOMET-lite achieves 0.363. Notably, the model quantized into 4-bit mode yields a slightly higher correlation for pairs with references, namely 0.379. Furthermore, quantization substantially reduces memory usage. The 8-bit quantization decreases the peak memory consumption of the distilled model by 17% from 1.8 GB to 1.5 GB, while the 4-bit quantization further reduces it to 1.4 GB. These results demonstrate that quantization is a viable option for further compressing the distilled model without substantial quality degradation. See Table 5 in Appendix D for full results.